# Exploring Model Kinship for Merging Large Language Models

### v1

@@ Abstract
Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs). However, our understanding of the expected performance gains and principles when merging any two models remains limited. In this work, we introduce model kinship, the degree of similarity or relatedness between LLMs, analogous to biological evolution. With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models. Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can yield better performance on benchmark datasets. Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
@@ 1 Introduction
Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023). However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang & Yang, 2022; Lu et al., 2024; Liu et al., 2024). Recent studies suggest that model merging (Singh & Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena & Raffel, 2022; Yang et al., 2024a) offers a viable approach for achieving multitask objectives by integrating multiple expert models. Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
To date, through model merging techniques, resercheres have developed many more powerful LLMs through iterative model merging (Beeching et al., 2023), and to some extent, achieved model evolution (Figure 1(c)). Despite these successes, progress has predominantly relied on trial and error, along with extensive human expertise, but lacks formalized guidance and standardized procedures. As the merging iterations progress, achieving further generalization gains becomes increasingly challenging (More details in Section 3). For example, as shown in Figure 1, model merging often resembles the process of hybrid evolution in biology, where the next generation may not show significant improvements or may even regress, highlighting the imperative for a deeper exploration of the underlying mechanisms driving these advancements.
To address this, we introduce model kinship, a metric inspired by the concept of kinship (Sahlins, 2013) from evolutionary biology (Figure 1(a)). This metric is designed to estimate the degree of similarity or relatedness between LLMs during the iterative model merging process, offering insights intended to enhance the effectiveness of the merging strategy. We utilize the model kinship to conduct a comprehensive analysis of model merging experiments from two perspectives: the overall merging process, including various independent merge experiments and the evolution path of specific models, demostrating the complete merging trajectory.
@@
Figure 1: An intuitive comparison between wheat evolution and model evolution. An interesting parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b). In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits. Similarly, in model merging, the merged model inherits parameters or weights from the contributing models. Part c demonstrates the iterative execution of model evolution. Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration. To be noted, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
Model kinship correlates with average performance gain in model merging. Emperically, we find that there is a strong correlation between variations in multitask capability, estimated by average task performance, and model kinship, which can help guide our selection of candidate models. We also notice that the model merging process has two stages: learning, where models experience significant performance improvements and saturation, where further improvements diminish and eventually stagnate. We think that the stagnation of improvements may be due to convergence in weight space, suggesting the presence of optimization challenges like local optima traps.
@@
Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship. Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping to mitigate degradation and avoid local optima during model evolution. Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process. Overall, this paper makes three key contributions:
@@
    Introduction of Model Kinship: We introduce model kinship, designed to assess the degree of similarity or relatedness between LLMs during the merging process, which can guide model merging strategies and holds promise for advancing auto-merging research.
@@
    Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging. Our findings highlight the dynamics of multitask performance improvement and stagnation. Additionally, we propose a preliminary explaination of the underlying mechanisms using model kinship.
@@
    Practical Model Merging Strategies using Model Kinship: We demostrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
@@ 2 Background
2.1 Model Merging
Model merging aims to integrate two or more domain-specific models into a unified framework, thereby harnessing their compositive capabilities across multiple tasks (Sung et al., 2023). While this approach shares conceptual similarities with ensemble methods (Dietterich et al., 2002; Dong et al., 2020; Jiang et al., 2023b), model merging generates a single, generalized model, avoiding the increased inference time associated with ensembles. Let $f_{i}$ represent the $i$-th model for merging, each with its unique parameters $\theta_{i}$. If the merging process follows method $\mathcal{F}$, the prediction $\hat{y}$ of the merged model $f_{\text{merge}}$ for input $x$ is:
@@ 2.1 Model Merging
When employing parameter averaging methods, the merged model preserves the same architecture and parameter size as the original models, enabling it to be reused in future merging processes [^2] This paper focuses solely on model merging with identical architectures, where the merged model retains the same parameters as before the merge.. This makes it feasible to iteratively enhance the model through repeated applications of model merging. In this paper, we refer to this process as “Model Evolution”. Empirical evidence on open LLM leaderboard (Beeching et al., 2023) demonstrates that model evolution can successfully produce highly generalized models, often outperforming their original ones (Maxime Labonne, 2024).
@@ 2.2 Model Kinship
Intuitively, understanding the task capabilities of models is crucial for effective model merging. However, the black-box nature of LLMs presents significant challenges in comprehending their advantages Shi et al. (2024) and task-specific capabilities. For fine-tuned models, task benchmark results (Gao et al., 2024; Li et al., 2023c) and the characteristics of the fine-tuning datasets are often used as indicators of task-related knowledge. Yet, these metrics often fall short in comprehensive comparisons. The internal relationships between tasks and datasets can arise confusion, particularly in complex merging scenarios that involve multiple tasks.
Motivated by the parallel drawn between biological reproduction and the process of model evolution (as seen in Figure 1), we propose that the concept of kinship, which is central to evolutionary biology for understanding breeding relationships and human genealogies (Thompson, 1985), can be conceptually extended to the field of model merging, specifically to describe the kinship between models. To be noted, Ilharco et al. (2023) has proposed Task Arithmetic and offered new insights into how the merging process preserves and blends task information from the original models. (Ilharco et al., 2023) propose to use task vectors, defined as the differences in weight space between fine-tuned models and a pre-trained model, and demonstrate that arithmetic operations on these task vectors can serve as an effective method for model merging.
Inspired by this idea, we introduce the concept of Model Kinship, a metric designed to assess the degree of similarity or relatedness between large language models (LLMs) based on their “genetic” information (a.k.a. the changes in weights during model evolution). Considering two target models involved in a single merge, the weights of each target model are denoted as $\theta_{\text{target}}\in\mathbb{R}^{d}$. In this paper, it is significant to claim that a target model may either be a fine-tuned model or a previously merged model. Similarly, $\theta_{\text{base}}\in\mathbb{R}^{d}$ represents the weights of the base model from which the two target models are derived. The difference of weights $\delta_{\text{target}}$ between the target model and the base model is denoted as:
$\delta_{\text{target}}=\theta_{\text{target}}-\theta_{\text{base}}.$
For a model merging process where models have been fine-tuned or merged from the same base model, the model kinship $r$ can be defined using the similarity metric function $sim(\cdot,\cdot)$, which measures the similarity between two model weight differences. For $n$ models, we formally define the model kinship $r$ as:
$r=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),$
where $\delta_{\text{i}}$ and $\delta_{\text{j}}$ are the weight differences between the $i$-th and $j$-th target models and the base model, respectively.
3 Preliminary Analysis of Model Kinship
We begin with a preliminary analysis of model kinship with open-sourced LLMs from the community to better understand model evolution.
@@ 3.1 Evaluation Metrics
3.2 Correlation Analysis of Model Kinship and Performance Gain
@@
In this analysis, we examine the distribution of merge gain and model kinship based on Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED) in open-sourced LLMs, originating from the Mistral-7B (Jiang et al., 2023a). Those models are obtained from the HuggingFace, with assistance from the Open LLM Leaderboard (Details in Appendix A.).
3.2.1 Results
Figure 2: Distribution of Sample Experiments: Relationship Between Model Kinship (X-axis) and Merge Gain (Y-axis). Model Kinships are calculated using the Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED).
Figure 2 illustrates the distribution of model kinship based on three similarity metrics (PCC, CS, ED) in relation to merge gain. The scatter plots reveal a moderate correlation between model kinship and merge gain, as indicated by the trend lines. To further quantify these relationships, the correlation value (use Pearson Correlation Coefficient) between model kinship and merge gain are calculated, as detailed in the second column of Table 1. While moderate correlations are observed for all three metrics, the corresponding p-values indicate a weak level of statistical significance, ranging from 0.05 to 0.1. In contrast, when examining absolute merge gain, we find stronger and statistically significant correlations, as shown in the third column of Table 1. These results suggest that model kinship alone is insufficient for predicting whether a model can achieve generalization gains through merging. However, it may serve as a key factor in determining the upper limit of merge gains, thus indicating potential improvements. Since no significant differences are observed among the three metrics, we will focus solely on model kinship based on PCC in the following sections to simplify the demonstration.
3.3 Sequence Analysis of Model Evolution Paths
In this analysis, we examine changes in performance and model kinship across independent model evolution paths to identify the phased pattern of the merging process. We focus on the yamshadow experiment 28-7B (Labonne, 2024), a Mistral 7B architecture model ranked as the top 7B merged model on the Open LLM Leaderboard. From its model family tree, we extract two primary merging paths: Path 1 and Path 2.
3.3.1 Results
Figure 3: Change in Average Task Performance and Merge Gain across the Model Evolution process: The selected paths originate from two distinct initial models, with the saturation stage observed after the vertical line. Note that the generation of Path 2 is aligned with Path 1 for demonstration purposes.
Figure 4: Comparision between Model Kinship (measured by Pearson Correlation Coefficient) and Average Task Performance (normalized to the same value scale).
We first focus on the average task performance and merge gains throughout the model evolution path (Figure 3.) Detailed data and branch information are summarized in Appendix A). Our observations indicate that the performance improvements of the iterative merging process are not linear and can be divided into two stages:
  • Learning Stage. In this stage, the average task performance generally experiences a rapid increase. Noticeable merge gains suggest that the merged models are continually acquiring multitask capabilities through the merging process.
  • Saturation Stage. As the process continues, improvements begin to plateau. During this stage, the merge gains approach zero, indicating that the model can no longer benefit from the merging process and has ceased to improve.
Additionally, we compare the trend of model kinship with average task performance. Figure 4 illustrates the changes in model kinship alongside average task performance (normalized to the same range as the corresponding metric) throughout the model evolution paths. In addition to the notable correlation observed between average task performance and model kinship throughout the evolution process, model kinship exhibits a similar stage-specific pattern, particularly evident in the saturation stage, suggesting a potential relationship with the underlying cause of saturation.
3.4 Analysis of the Model Kinship in Different Merging Stages
Findings in previous analysis raises the question of “how model kinship affects performance improvements in model evolution”. To investigate the causality between model kinship and the stagnation of improvements, we examine the variation of model kinship across different merging stages from a broader perspective.
Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of top-performing models at each stage. For this analysis, we randomly select 5 models from each merging stage - Saturation Stage ($\geq 0.75$), Learning Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
Figure 5 illustrates the model kinship between models within each group. Our analysis reveals that model kinship increases with the average task performance, even across models that follow different evolution paths yet share similar average performance levels. Additionally, during the saturation stage, all potential merges display a strong affinity, with model kinship values nearing 1. Since model kinship indicates the similarity of weights, we concluide the final findings as:
Findings: Model merging experiences a saturation stage, where the model kinship among top-performing models increases throughout the iterative merging process. This implies that the models converge to similar forms, resulting in excessive relatedness that undermines the effectiveness of the model merging strategy.
4 Using Model Kinship to Improve Model Merging
Inspired by the above findings, we further leverage model kinship to enhance the model merging process. We firstly conduct experiments employing a performance-prior greedy merging strategy. Note that the greedy strategy may eventually lead to convergence, we further introduce Top-$k$ Greedy Merging with Model Kinship (Algorithm 1). Our results indicate that while the greedy strategy focuses on short-term gains, it can lead to parameter convergence and suboptimal outcomes. By integrating model kinship, we can help the strategy avoid local optima. Furthermore, we find that model kinship holds potential for enhancing merging strategies as an early stopping criterion.
4.1 Experiment Setup
Algorithm 1 Top $k$ Greedy Merging with Model Kinship.
1:  Generate the first generation of merged models by merging each pair in set $M$$\{M_{1},M_{2},\dots,M_{n}\}$.
2:  Evaluate each merged model using $f$ and select the top $k$ models. Denote this set as $S=\{M_{1},M_{2},\dots,M_{n}\}$.
3:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $m$ models from the previous iteration.
4:  while $S\neq S_{\text{prev}}$ do
5:     Set $M_{\text{prev}}=M$.
6:     Set $S_{\text{prev}}=S$.
7:     Select $k$ pairs of models from $S$ with the highest performance according to $f$.
8:     Identify the current best model $M_{best}\in S$.
9:     Identify the model $M_{f}\in S$ with the highest model kinship to $M_{best}$ from the $M_{\text{prev}}$ according to the similarity metric $sim(\cdot,\cdot)$.
10:     Merge $M_{f}$ with $M_{best}$ to generate a new model $M_{\text{exp}}$ and add $M_{\text{exp}}$ into set $M$.
11:     Merge each selected pair to $M_{merged}$ (named as Model-gen-id) for merged models and add merged models into set $M$.
12:     Evaluate each new model using $f$ and update $S$ to be the new top $k$ models.
13:  end while
Note: The blue-highlighted steps are only executed in modified experiments incorporating model kinship-based exploration.
Figure 6: Left (a): The comparison of task performance improvement across merging generations. The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities. Right (b): The partial model family tree from the controled experiments. The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
LLMs.
We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210. Each model demonstrates distinct strengths across three benchmarks: Winogrande (5-shot), GSM8k (5-shot), and TruthfulQA MC2 (0-shot). All models are evaluated using the Language Model Evaluation Harness (Gao et al., 2024), a widely adopted framework for testing LLMs.
Merging Method.
We conduct two iterative model merging experiments, both utilizing the SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for merging [^3] We provide a Google Colab Notebook.. For implementation, we employ Mergekit (Goddard et al., 2024), a comprehensive toolkit that offers simple access to state-of-the-art model merging techniques.
Top $k$ Greedy Merging.
This strategy utilizes the vanilla Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1). This approach has been widely adopted in the community and has demonstrated notable success. In Figure 6 (b), models generated by the greedy strategy are indicated in green, while the best-performing models are highlighted in red.
Top $k$ Greedy Merging with Model Kinship.
The propposed strategy simply introduces an additional exploration step, based on model kinship, to the original greedy strategy (highlighted by the blue part in Algorithm 1). This approach aims to merge the best-performing model with the model that has the most distinct task capabilities, in order to discover potentially better solutions. In Figure 6 (b), models generated by our strategy are marked in purple, while the best-performing models are marked in yellow.
4.2 Results and Discussion
Figure 6 (a) illustrates the improvements in top average task performance across merging generations. Table 2 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
Both strategies achieve the multitask goals. However, the vanilla greedy strategy stops improving after Generation 2, stabilizing at an average task performance of 68.72. In contrast, Experiment 2, utilizing model kinship-based exploration, escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
Table 2: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach. The first three models serve as the foundation models in both experiments. Merged models are labeled using the structure: Model-{generation number}-{model identification number}.

### v2

@@ Abstract
Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs). The open-source community has driven model evolution by iteratively merging existing models. However, a principled understanding of the expected gains and underlying factors in model merging remains lacking. In this work, we examine model evolution through continual merging, analogous to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs. With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models. Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can yield better performance on benchmark datasets. Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
Exploring Model Kinship for Merging Large Language Models
Yedi Hu, Yunzhi Yao, Shumin Deng , Huajun Chen, Ningyu Zhangthanks: Corresponding Author. Zhejiang University   National University of Singapore, NUS-NCS Joint Lab, Singapore {zhangningyu}@zju.edu.cn
@@ 1 Introduction
Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023). However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang and Yang, 2022; Lu et al., 2024; Liu et al., 2024).
@@
Figure 1: An intuitive comparison between wheat evolution and model evolution. An interesting parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b). In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits. Similarly, in model merging, the merged model inherits parameters or weights from the contributing models. Part c demonstrates the iterative execution of model evolution. Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration. Notably, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multitask objectives by integrating multiple expert models. Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
To date, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023). Many of these models are created through a biologically inspired evolutionary process that involves iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)). Despite these successes, the current approach faces critical limitations. Progress often relies on trial and error and extensive human expertise, with little formal guidance or standardized procedure. To address this problem, we propose a strategic model evolution framework (Figure 1(c)) that leverages explicit strategies to guide the direction of model evolution toward improved performance. We show that even a simple greedy strategy can outperform baseline merging approaches.
However, in the late stages of both the community experiments and the greedy strategy evolution, achieving further gains in generalization becomes increasingly difficult. To explore possible solutions, we introduce ’model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process. Model kinship is designed to quantify the degree of similarity or relationship between models throughout the iterative merging process. By providing a principled way to measure these relationships, model kinship offers crucial insights for refining merging strategies.
We conduct a comprehensive analysis of model merging experiments based on model kinship. We observe that the community model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau. Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of the potential effectiveness of a merge. These findings inspire two main insights: (1) high-kinship merges can lead to performance stagnation, akin to inbreeding; (2) low-kinship merges carry greater risk but may yield larger gains and facilitate escape from local optima.
@@
Inspired by this, we propose a new continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship.  [^2] https://github.com/zjunlp/ModelKinship.. Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping mitigate degradation and avoid local optima during model evolution. Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process.
In general, this paper makes four key contributions:
@@
    Continual Model Merging as a Feasible Framework for Model Evolution: We propose continual model merging as a viable framework for evolving large language models. Through strategically guided merging across iterations, this approach yields consistent improvements in generalization and task performance.
@@
    Introducing Model Kinship: We introduce model kinship, designed to assess the degree of similarity or relationship between LLMs during the merging process, which can guide model merging strategies and holds promise for advancing auto merging research.
@@
    Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging. Our findings highlight the dynamics of multitask performance improvement and stagnation during evolution. In addition, we propose a preliminary explanation of the underlying mechanisms using model kinship.
  4. 4.
    Practical Model Merging Strategies using Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
@@ 2 Background
2.1 Model Merging: Fundamentals
Model merging aims to integrate two or more domain-specific models into a unified framework, thereby harnessing their composite capabilities across multiple tasks (Sung et al., 2023). While this approach shares conceptual similarities with ensemble methods (Dietterich et al., 2002; Dong et al., 2020; Jiang et al., 2023b), model merging generates a single, generalized model, avoiding the increased inference time associated with ensembles. Let $f_{i}$ represent the $i$-th model for merging, each with its unique parameters $\theta_{i}$. If the merging process follows method $\mathcal{F}$, the prediction $\hat{y}$ of the merged model $f_{\text{merge}}$ for input $x$ is:
@@ 2.1 Model Merging: Fundamentals
2.2 Continual Model Merging: Benefits and Challenges
Parameter averaging methods enable the merged model to retain the same architecture and parameter size as the original models, allowing for reuse in future merging processes. Leveraging this property, the community has progressively enhanced models through repeated applications of model merging, a process we refer to as “Model Evolution”. Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that model evolution can yield highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
However, current model evolution practices typically involve random merging by multiple contributors, leading to high computational costs and unstable behavior that limit its feasibility for industrial applications.
3 Strategic Model Evolution
In this section, we demonstrate how strategic model evolution via continual merging can achieve better generalization across tasks.
3.1 Method
We first define an iterative process for model evolution, where models are incrementally merged to produce a more generalized model. The direction of model evolution is guided by a principled strategy that governs the model selection. At each iteration $t$, a selection strategy $\mathcal{S}$ is applied to the current pool of models $\mathcal{P}_{t}=\{M_{t}^{(1)},M_{t}^{(2)},\ldots,M_{t}^{(k)}\}$, producing a subset of candidate models $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$. These selected models are then combined using a merge method $\mathcal{M}$, such as weighted averaging or linear interpolation in the weight space, resulting in the next-generation model $M_{t+1}$. This iterative process continues until a predefined end strategy $\mathcal{E}$ is matched. Formally, the process is defined as:
$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{subject to}\quad\neg\mathcal{E}(M_{t+1})$ (2)
Method TruthfulQA Winogrande GSM8k Avg.
Ties 62.76 78.61 0.11 47.16
Linear 56.37 78.08 68.54 67.66
Continual Merge (ord1) 47.15 76.24 53.15 58.84
Continual Merge (ord2) 61.01 79.56 63.76 68.11
Continual Merge (ord3) 49.80 78.53 55.72 61.35
Strategic Model Evolution 50.94 80.11 75.13 68.72
Table 1: Performance Comparison across different merging methods.
3.2 Setup
Baseline Method.
We adopt two commonly used approaches for merging multiple models as baselines. The first is multi-model merging, where all models are merged simultaneously. The second is sequential continual merging, where three models are merged in sequence using a pairwise merging strategy. For these baselines, we apply TIES (Yadav et al., 2023) and Linear Merge for the multi-model merging case and SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for the pairwise merging in the sequential strategy.
Top $k$ Greedy Merging.
Our method applies continual model merging with greedy strategy, the Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1 in Appendix C). Each single merging step in the continual merging experiments is performed using SLERP.
Models and Datasets.
We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
Evaluation.
Evaluation is performed using three task-specific benchmark datasets: Winogrande, GSM8k, and TruthfulQA. These benchmarks demonstrate the distinct strengths of the three selected fine-tuned models. Further details on the tasks are provided in Appendix A.3.
3.3 Results
As shown in Table 1, Strategic Model Evolution can yield better generalization when combining multiple tasks. However, compared to simple sequential merging, the improvements are limited due to the simplicity of the greedy strategy. This observation raises an important question regarding optimal merging strategies:
Problem: How can we design more principled and effective strategies for model merging to better facilitate model evolution?
4 Preliminary Analysis of Model Kinship
To explore efficient strategies for model evolution, we propose a novel metric, model kinship, which captures task-related differences between models and may help predict merging outcomes. We then conduct a preliminary analysis of model kinship in community merging experiments on LLMs to evaluate its potential for informing and enhancing model evolution strategies.
@@ 4.1 Model Kinship
Drawing inspiration from the parallel between artificial selection and model evolution (as detailed in Appendix G), we hypothesize that a concept analogous to kinship, the genetic relatedness studied in evolutionary biology (Thompson, 1985), can also apply to model merging. Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between models. This analogy suggests that just as genetic kinship influences breeding outcomes, model kinship may similarly impact the effectiveness of merging strategies in enhancing generalization performance.
We adopt the most intuitive representation, inspired by the cosine similarity analysis introduced in the Task Arithmetic paper (Ilharco et al., 2023). This metric is designed to evaluate the degree of similarity or relatedness between the task capabilities of large language models (LLMs) based solely on their "genetic" information, meaning the changes in their weights, during model evolution. Considering two models $m_{i}$, $m_{j}$ involved in a model evolution originated from the pre-trained model $m_{base}$, the weights of $m_{i}$, $m_{j}$ are denoted as $\theta_{\text{i}},\theta_{\text{j}}\in\mathbb{R}^{d}$. Similarly, $\theta_{\text{base}}\in\mathbb{R}^{d}$ represents the weights of the pre-trained model. Since the differences between models emerge after fine-tuning and merging, the variation of weights during model evolution is crucial. It is calculated as:
$\delta_{\text{i}}=\theta_{\text{i}}-\theta_{\text{base}},\delta_{\text{j}}=\theta_{\text{j}}-\theta_{\text{base}}$ (3)
Model kinship $r$ is designed to capture the similarity of task capabilities between models. In this paper, we explore multiple potential metrics for evaluating similarity. For the calculation, $sim(\cdot,\cdot)$ denotes the similarity metric function used. Considering two cases merging of 2 models and merging of $n$ models, we formally define model kinship $r$ as:
$r=\begin{cases}sim(\delta_{1},\delta_{2}),&\text{(Merge 2)}\\ \frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),&\text{(Merge N)}\end{cases}$ (4)
We investigate the relationship between task performance and model kinship (see Appendix E for the full analysis). The results reveal strong correlations, reinforcing the view that model kinship reflects task-related differences between models.
@@ 4.2 Evaluation Metrics
Figure 2: Distribution of Sample Experiments: Relationship Between Model Kinship (X-axis) and Merge Gain (Y-axis). Model kinships are calculated using PCC, CS, and ED.
Figure 3: Change in Average Task Performance and Merge Gain across the Model Evolution Process: Paths originate from two different base models. The vertical line marks the transition to the saturation stage. Path 2 is temporally aligned with Path 1 for clarity.
4.3 Analysis of Model Kinship: Correlation and Evolution Dynamics
In this section, we analyze model kinship from two perspectives: (1) its correlation with performance gain across a broad range of open-sourced LLM merges, and (2) its dynamic along specific model evolution paths. These analyses aim to clarify the relationship between model kinship and multitask capability improvements, as well as to identify phases of merge effectiveness.
4.3.1 Correlation Between Model Kinship and Performance Gain
We begin by exploring the potential relationship between merge gain and model kinship using three similarity metrics: Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED). The models used in this analysis are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
As illustrated in Figure 2, the scatter plots derived from all three metrics suggest a moderate correlation between model kinship and merge gain. Table 2 reports Pearson correlation values for both signed and absolute merge gains. While the correlations for signed gains appear relatively weak (with p-values between 0.05 and 0.1), those for absolute merge gains are comparatively stronger and show greater statistical significance. These observations imply that model kinship may offer some indication of the potential magnitude of merge gains, though it appears less effective at predicting the direction of change. While we cannot assert a causal relationship, the association provides useful insight into how kinship might relate to merge outcomes. In light of the comparable performance across the three metrics, we use PCC-based kinship in the remainder of our analysis for consistency.
4.3.2 Model Kinship in Evolution Paths
As a further exploration, we examine model kinship across independent model evolution paths to investigate potential phase patterns in the merging process. This analysis centers on the yamshadow experiment 28-7B (Labonne, 2024), a Mistral-7B-based model that ranks among the top-performing merged models on the Open LLM Leaderboard. From its model family tree, we extract two main merging trajectories, referred to as Path 1 and Path 2, for comparison.
Figure 4: Comparison between Model Kinship (PCC) and Average Task Performance (normalized to the same scale).
Figure 5: Left (a): The comparison of task performance improvement across merging generations. The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities. Right (b): The partial model family tree from the controlled experiments. The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
Figure 3 displays the average task performance and the merge gains along the two evolution paths. The merging process exhibits two distinct phases:
  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multitask balance.
  • Saturation Stage. Performance stabilizes, and additional merges result in minimal or no measurable improvement.
Figure 4 shows how the model kinship and normalized average performance change over the course of the merging process. Both metrics display a similar two-stage trend: they increase during the Improving Stage and level off during the Saturation Stage. This parallel trajectory suggests a potential link between model kinship and performance gains, indicating that improvements in generalization may coincide with and possibly depend on increases in model kinship.
@@ 4.3.2 Model Kinship in Evolution Paths
Additionally, to generalize the findings from the Evolution Paths, we analyze how model kinship evolves across different stages of the merging process, thereby broadening the scope from individual paths to the full evolution picture. (detailed in Appendix D.1). The results show that the best performing models exhibit high mutual kinship, which may lead to a stagnation stage in the merging process.
4.4 Discussion
Considering all results that we observed, this analysis provides two main inspirations for the application of model kinship:
  • High kinship merges may lead to performance stagnation, similar to biological inbreeding.
  • Low kinship merges involve risk, but can lead to greater gains, potentially enabling escape from local optima caused by greedy strategy.